a vertical cyclinder of height 100cm contains air at a constant tempreature.the top is closed by a frictionless light piston.the atmospheric pressure is equal to 75cm of mercury.mercury is slowly poured over the piston.find the maximum height of the mercury column that can be put on yhe piston.pls also explain the concept behind this i m not able to such type of problems in which v have to calculate atmospheric pressure+ghdensity?             




 

 


  

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

let x be maximum height of mercury column. P1V1=P2V2, so pressure of air inside cylinder=(100/(100-x))*75=75+x,

upon solving this equation for x, you get the answer